Show the cash flows for the following four bonds, each of which has a par value...

  1. Show the cash flows for the following four bonds, each of which has a par value of $1000 and pays interest semiannually.

Bond

Coupon (%)

Maturity (years)

Price ($)

W

7

5

884.20

X

8

7

948.90

Y

9

4

967.70

Z

10

10

456.39

  1. Calculate the yield to maturity for the four bonds.  
Answer #1

YTM = Interest + (Redemption Value - Market Value/Years to maturity (Redemption Value + Market Price)/2 [35+(35*1.035)]+(1000[45+(45*1.045)]+(1000-967.70)/4 (1000+967.70)/2 Bond Y= 92.025+8.075 983.85 =9.99% [50+(50*1.05)]+(1000-456.39)/10 (1000+456.